# Yong Liu    Guo Qin    Zhiyuan Shi    Zhi Chen    Caiyin Yang    Xiangdong Huang    Jianmin Wang    Mingsheng Long

### v2

@@ 1 Introduction
Time series forecasting has fascinated people for thousands of years. Although people have been able to determine the time using instruments like sundials in 3000 BC, time series forecasting is intrinsically non-deterministic (Box et al., 2015). Therefore, generating a variety of probable predictions is crucial for decision-making. The growing demand has facilitated numerous statistical approaches over the past decades (Hyndman, 2018; Box, 2013), which provide high-profile theories and probabilistic tools for making reliable schedules. Recent advancements bring the boom of deftly designed models that automatically learn intricate dynamics and correlations from raw data (Oreshkin et al., 2019; Nie et al., 2022; Zhang & Yan, 2023; Liu et al., 2023a). Despite the impressive performance, deep models necessitate task-specific training on sufficient in-distribution data. Motivated by advances in large models (Bommasani et al., 2021), pre-trained time series foundation models have shown promising capabilities in out-of-distribution tasks (Das et al., 2023b; Liu et al., 2024b; Woo et al., 2024; Ansari et al., 2024).
@@
Current research on time series foundation models has converged on building unified, scalable, and out-of-the-box forecasters, exhibiting zero-shot performance close to or sometimes surpassing supervised methods (Aksu et al., 2024). Notably, Transformers (Radford et al., 2018) are currently the de facto architecture of these models. While pre-trained Transformers with an inherent generative ability have facilitated great success in language, image, and video generation (Ramesh et al., 2021; OpenAI, 2023; Liu et al., 2024c), most time series foundation models are not “generative” or, more specifically, probabilistic forecasters, thereby limiting reliability in decision-making. Although parametric densities specified with prior distributions (Wen et al., 2017; Woo et al., 2024) can be adopted to address uncertainty in time series forecasting, they can reduce the capacity of distributions learned by pre-trained models, especially on time series modality characterized by high heterogeneity. To learn arbitrarily intricate distributions without mode collapse, language modeling (Bengio et al., 2000) that learns the categorical distribution via cross-entropy loss inspires subsequent works (Gruver et al., 2023; Ansari et al., 2024), which treat time series as a foreign language using discrete tokenization. Still, discrepancies between continuous-valued time series and discrete language tokens can lead to out-of-vocabulary issues and coarse-grained prediction intervals.
@@
As shown in Figure 1, Sundial is presented as the first family of generative models among time series foundation models. As foundation models intend to learn complicated distributions from extensive datasets and facilitate transferability across agnostic downstream datasets, we do not specify any prior parametric densities, such as unimodal and multimodal Gaussian mixtures. Instead, we delve into generative modeling to tame Transformers as native, flexible, and scalable time series foundation models. By comparing to denoising diffusion models (Li et al., 2024), we opt for a simple yet effective flow-matching framework (Lipman et al., 2022), which provides notable efficiency and sample quality (Tong et al., 2023). We propose TimeFlow Loss, a parameterized training objective (Zhang et al., 2018) for autoregressive models to learn and sample from each token’s predictive distribution. Optimizing models in the original continuous-valued domain, TimeFlow Loss facilitates patch-level generation and enables fast inference, which is naturally compatible with the time series modality.
@@
In addition to TimeFlow, we enhance the Transformer with minimal but critical adaptations. We develop feasible patch tokenization for arbitrary-length input time series. We adopt RoPE (Su et al., 2024), Pre-LN (Xiong et al., 2020), FlashAttention (Dao et al., 2022), and KV Cache (Pope et al., 2023), which are crucial but generally neglected in the development of time series foundation models. Besides, we pre-train our models by multi-patch prediction to reduce autoregression steps. We realize a rapid generation of multiple samples by reusing a shared lookback representation. Beyond facilitating scalable pre-training, these adaptations help real-time long-context inference and long-term generation.
@@
To validate the scaling law of time series foundation models, we collect and curate TimeBench with an unprecedented volume of a trillion time points. We present Sundial as a family of highly capable foundation models, which achieve state-of-the-art on three large-scale and best-recognized benchmarks, including Time-Series-Library (TSLib) (Wu et al., 2022), GIFT-Eval (Aksu et al., 2024), and FEV (Ansari et al., 2024). Our contributions lie in these aspects:
@@ 2.1 Time Series Forecasting
Forecasting is essential for decision-making. Advancements in deep learning for time series include theory-inspired deep modules (Wu et al., 2021; Liu et al., 2023b; Wu et al., 2022), architecture-oriented adaptations (Bai et al., 2018; Salinas et al., 2020; Lim et al., 2021), and time series preprocessing (Kim et al., 2021; Nie et al., 2022). Deep models learn the dataset-level distribution and benefit from strong generalization and model capacity. Statistical methods conduct case-by-case fitting on input series, achieving notable performance on small data (Ke et al., 2017; Hyndman, 2018).
One of the efforts towards more capable forecasters focuses on the foundation models (Bommasani et al., 2021), which address data-scarce scenarios by pre-training. More capable models support zero-shot forecasting, making inferences as fast as statistical methods and possessing large model capacity as deep models. Another aspect is to address uncertainty in time series forecasting. There is a growing research emphasis on probabilistic forecasting (Woo et al., 2024; Ansari et al., 2024). While parametric densities can be adopted as training objectives of probabilistic forecasting, they can be too specific to meet the heterogeneity of large-scale datasets, resulting in mode collapse in representation learning and over-smooth predictions (Figure 14-15). In this work, we introduce generative time series foundation models, which naturally address the uncertainty in forecasting.
@@ 2.2 Time Series Foundation Models
Recent research has concentrated on building versatile large time series models (Liang et al., 2024). With the advances made in large language models, Transformer has become the dominant architecture. Several works adapt Transformers to address the unique 2D-dimensionality and heterogeneity of time series (Woo et al., 2024; Liu et al., 2024a). Specifically, our work delves into tokenization and optimization. Models such as TimesFM (Das et al., 2023b), Timer (Liu et al., 2024a, b), and Time-MoE (Shi et al., 2024b) embed continuous values and fit unimodal distributions via MSE or quantile loss (Wen et al., 2017). However, prior loss may result in mode collapse because predictive distributions are highly divergent across different domains. Besides, these models cannot provide the confidence level of predictions, limiting reliability for decision-making. Based on continuous tokenization, Moirai (Woo et al., 2024) presents a probabilistic model learning a mixture of distributions, but this prior can still fail to accommodate complex distributions. Inspired by language modeling, Chronos (Ansari et al., 2024) discretizes series via bucket quantization, learning more flexible categorical distributions by cross-entropy. Still, discrete tokenization is applied at each time point, which can lead to long contexts. Also, the final performance can be sensitive to quantization techniques. Unlike before, we tame Transformers as native time series foundation models, learning flexible distributions without discrete tokenization.
@@ 2.3 Generative Modeling for Time Series
By addressing complicated distributions during pre-training, generative modeling has become a focal point in the development of various foundation models (Zhao et al., 2023; Liu et al., 2024c). While this direction for time series mostly focused on time series generation (Tashiro et al., 2021) and task-specific forecasters (Rasul et al., 2021; Shen & Kwok, 2023; Kollovieh et al., 2024), generative modeling for time series foundation models is hardly explored. With the comparable flexibility in distribution learning as language modeling, diffusion denoising (Sohl-Dickstein et al., 2015) and flow-matching (Lipman et al., 2022) have gained increasing prevalence in continuous-valued modalities (Lipman et al., 2024). Compared with diffusion denoising models, flow-matching provides a simple yet efficient framework. With fewer steps involved in the forward and reverse processes, large models based on flow-matching have shown superior performance in image generation (Esser et al., 2024).
Despite the connection in value continuity, generating images and future time series are fundamentally different tasks due to the autoregressive property of forecasting. Our proposed TimeFlow Loss is designed for autoregressive models to conduct conditional generation, which is a parameterized loss function (Zhang et al., 2018) for arbitrary distributions and enhances representation learning of foundation models.
@@ 3.1 Flow-Matching
Furthermore, Lipman et al. (2022) proved the equivalence of optimizing the Conditional Flow-Matching objective:
@@ 3.2 Generative Models for Probabilistic Forecasting
Given a historical observation $x_{1:t}=\{x_{1},...,x_{t}\}$, the target of time series forecasting is to predict future time series $x_{t+1:t+f}=\{x_{t+1},\dots,x_{t+f}\}$. The task can be generally formulated as $p(x_{t+1:t+f}|\mathbf{h}_{t})$, where $\mathbf{h}_{t}=f_{\phi}(x_{1:t})$ is the learned representation from a deep model $f_{\phi}$. In probabilistic forecasting, explicit optimization objectives are utilized to predict the statistics of future series, e.g., MSE or quantile loss, which have specified $p$ as a prior distribution. While using one parametric density generally fits well on a small amount of data, it can be the major bottleneck for scaling time series foundation models. Inspired by the success of large generative models (Rombach et al., 2022; OpenAI, 2023; Esser et al., 2024), we introduce generative modeling to realize probabilistic forecasting:
$p(x_{t+1:t+f}|\mathbf{h}_{t})=g_{\theta}\big{(}f_{\phi}({x_{1:t}})\big{)}.$ (3)
$g_{\theta}$ is a small trainable generative network conditioned on the learned representations of $f_{\phi}$, which is jointly optimized with $f_{\phi}$. While the generative model automatically fits the target distribution, it can sample raw predictions and calculating their statistics for probabilistic forecasting. The aim is conceptually related to conformal prediction (Vovk et al., 2005) but models uncertainty beyond prediction intervals.
@@ 4 Approach
In this work, we conduct a univariate pre-training paradigm, which adopts the S3 format proposed by Liu et al. (2024b) to address multivariate data. To mitigate value range discrepancy, we conduct normalization on time series individually per variable. Afterwards, we sample varying-length training samples with the maximum context length of $2880$. As a foundation model, Sundial is required to predict on out-of-distribution series with varied lengths during inference.
@@ 4.1.1 Time Series Tokenization
We adopt stationarization (Liu et al., 2022), a non-parametric two-stage instance normalization conducted within each sample, which is initially proposed to mitigate non-stationarity of time series. Here, it helps to address temporal distribution shift and outlier ranges in input series, improving generalizability for zero-shot forecasting.
@@
Given a univariate time series $\mathbf{X}=\{x_{1},\dots,x_{T}\}$, it is divided into patches $\mathbf{x}_{i}=x_{1+(i-1)P:iP}$ with the length of $P$. To address non-divisible length, we pad the input at the beginning and use a binary mask $\mathbf{m}_{i}\in\mathbb{R}^{P}$ for each patch to indicate the padded position. It will lead to $N=\lceil T/P\rceil$ such input tokens. Subsequently, we use a shared MLP $:\mathbb{R}^{2P}\mapsto\mathbb{R}^{D}$ to embed all patch tokens:
@@
where $\mathbf{h}_{i}\in\mathbb{R}^{D}$ and $D$ is the dimension of token embedding. Unlike point-level quantization (Ansari et al., 2024), we reserve original values without discrete quantization. It also reduces the context length (in token) of the Transformer.
@@ 4.1.2 Transformer Backbone
Given $N$ token embeddings $\{\mathbf{h}_{i}\}$, we adopt several crucial adaptations on a decoder-only Transformer to obtain per-token representations aggregated from all previous tokens. First, we adapt Pre-LN (Xiong et al., 2020) to improve pre-training stability. Second, we leverage a causal self-attention mechanism with RoPE (Su et al., 2024) that introduces the position information of patch tokens. It can be formulated as follows (the layer index is omitted for simplicity):
@@
where $\mathbf{W}_{\mathbf{q}},\mathbf{W}_{\mathbf{k}},\mathbf{W}_{\mathbf{v}}\in\mathbb{R}^{D\times d}$ project token embeddings $\mathbf{H}=\{\mathbf{h}_{i}\}$ into $d$-dimensional queries, keys, and values. $\mathbf{R}_{\Theta,t}\in\mathbb{R}^{d\times d}$ is the rotary matrix with rotation degree $(t\cdot\Theta)$. Lastly, we implement FlashAttention (Dao et al., 2022) and KV Cache (Pope et al., 2023), since these enhancements for deployment are increasingly emphasized in large foundation models (Shoeybi et al., 2019; Rasley et al., 2020).
@@ 4.1.3 TimeFlow Loss
Given representations $\{\mathbf{h}_{i}\}$ extracted by the last layer of the Transformer, we aim to generate length-$F$ predictions $\widehat{\mathbf{y}}_{i}=\widehat{x}_{1+iP,F+iP}$ at each position $i$ via our autoregressive model. Motivated by the empirical observation that a larger patch size improves the performance in decoder-only Transformers (Das et al., 2023b) while a small patch size can be more flexible to accommodate data of different frequencies, we adopt multi-patch predictions ($F>P$) for pre-training, which also reduces the steps of autoregressive inference.
@@
The training process involves sampling the noised $\mathbf{y}_{i}^{(t)}$, and jointly input it with $t$. The condition $\mathbf{h}_{i}$ is integrated into the flow-matching network via AdaLN (Peebles & Xie, 2023). TimeFlow Loss for autoregressive models is formulated as:

### v3

@@ 1 Introduction
Time series forecasting has fascinated people for thousands of years. Although people have been able to determine the time using instruments like sundials in 3000 BC, time series forecasting is intrinsically non-deterministic (box2015time). Therefore, generating a variety of probable predictions is crucial for decision-making. The growing demand has facilitated numerous statistical approaches over the past decades (hyndman2018forecasting; box2013box), which provide high-profile theories and probabilistic tools for making reliable schedules. Recent advancements bring the boom of deftly designed models that automatically learn intricate dynamics and correlations from raw data (oreshkin2019n; nie2022time; zhang2023crossformer; liu2023itransformer). Despite the impressive performance, deep models necessitate task-specific training on sufficient in-distribution data. Motivated by advances in large models (bommasani2021opportunities), pre-trained time series foundation models have shown promising capabilities in out-of-distribution tasks (das2023decoder; liutimer; woo2024unified; ansari2024chronos).
@@
Current research on time series foundation models has converged on building unified, scalable, and out-of-the-box forecasters, exhibiting zero-shot performance close to or sometimes surpassing supervised methods (aksugift). Notably, Transformers (radford2018improving) are currently the de facto architecture of these models. While pre-trained Transformers with an inherent generative ability have facilitated great success in language, image, and video generation (ramesh2021zero; openai2023gpt; liu2024sora), most time series foundation models are not “generative” or, more specifically, probabilistic forecasters, thereby limiting reliability in decision-making. Although parametric densities specified with prior distributions (wen2017multi; woo2024unified) can be adopted to address uncertainty in time series forecasting, they can reduce the capacity of distributions learned by pre-trained models, especially on time series modality characterized by high heterogeneity. To learn arbitrarily intricate distributions without mode collapse, language modeling (bengio2000neural) that learns the categorical distribution via cross-entropy loss inspires subsequent works (gruver2023large; ansari2024chronos), which treat time series as a foreign language using discrete tokenization. Still, discrepancies between continuous-valued time series and discrete language tokens can lead to out-of-vocabulary issues and coarse-grained prediction intervals.
@@
As shown in Figure 1, Sundial is presented as the first family of generative models among time series foundation models. As foundation models intend to learn complicated distributions from extensive datasets and facilitate transferability across agnostic downstream datasets, we do not specify any prior parametric densities, such as unimodal and multimodal Gaussian mixtures. Instead, we delve into generative modeling to tame Transformers as native, flexible, and scalable time series foundation models. By comparing to denoising diffusion models (li2024autoregressive), we opt for a simple yet effective flow-matching framework (lipman2022flow), which provides notable efficiency and sample quality (tong2023improving). We propose TimeFlow Loss, a parameterized training objective (zhang2018unreasonable) for autoregressive models to learn and sample from each token’s predictive distribution. Optimizing models in the original continuous-valued domain, TimeFlow Loss facilitates patch-level generation and enables fast inference, which is naturally compatible with the time series modality.
@@
In addition to TimeFlow, we enhance the Transformer with minimal but critical adaptations. We develop feasible patch tokenization for arbitrary-length input time series. We adopt RoPE (su2024roformer), Pre-LN (xiong2020layer), FlashAttention (dao2022flashattention), and KV Cache (pope2023efficiently), which are crucial but generally neglected in the development of time series foundation models. Besides, we pre-train our models by multi-patch prediction to reduce autoregression steps. We realize a rapid generation of multiple samples by reusing a shared lookback representation. Beyond facilitating scalable pre-training, these adaptations help real-time long-context inference and long-term generation.
@@
To validate the scaling law of time series foundation models, we collect and curate TimeBench with an unprecedented volume of a trillion time points. We present Sundial as a family of highly capable foundation models, which achieve state-of-the-art on three large-scale and best-recognized benchmarks, including Time-Series-Library (TSLib) (wu2022timesnet), GIFT-Eval (aksugift), and FEV (ansari2024chronos). Our contributions lie in these aspects:
@@ 2.1 Time Series Forecasting
Forecasting is essential for decision-making. Advancements in deep learning for time series include theory-inspired deep modules (wu2021autoformer; liu2023koopa; wu2022timesnet), architecture-oriented adaptations (bai2018empirical; salinas2020deepar; lim2021temporal), and time series preprocessing (kim2021reversible; nie2022time). Deep models learn the dataset-level distribution and benefit from strong generalization and model capacity. Statistical methods conduct case-by-case fitting on input series, achieving notable performance on small data (ke2017lightgbm; hyndman2018forecasting).
One of the efforts towards more capable forecasters focuses on the foundation models (bommasani2021opportunities), which address data-scarce scenarios by pre-training. More capable models support zero-shot forecasting, making inferences as fast as statistical methods and possessing large model capacity as deep models. Another aspect is to address uncertainty in time series forecasting. There is a growing research emphasis on probabilistic forecasting (woo2024unified; ansari2024chronos). While parametric densities can be adopted as training objectives of probabilistic forecasting, they can be too specific to meet the heterogeneity of large-scale datasets, resulting in mode collapse in representation learning and over-smooth predictions (Figure 14-15). In this work, we introduce generative time series foundation models, which naturally address the uncertainty in forecasting.
@@ 2.2 Time Series Foundation Models
Recent research has concentrated on building versatile large time series models (liang2024foundation). With the advances made in large language models, Transformer has become the dominant architecture. Several works adapt Transformers to address the unique 2D-dimensionality and heterogeneity of time series (woo2024unified; liu2024timer). Specifically, our work delves into tokenization and optimization. Models such as TimesFM (das2023decoder), Timer (liu2024timer; liutimer), and Time-MoE (shi2024time) embed continuous values and fit unimodal distributions via MSE or quantile loss (wen2017multi). However, prior loss may result in mode collapse because predictive distributions are highly divergent across different domains. Besides, these models cannot provide the confidence level of predictions, limiting reliability for decision-making. Based on continuous tokenization, Moirai (woo2024unified) presents a probabilistic model learning a mixture of distributions, but this prior can still fail to accommodate complex distributions. Inspired by language modeling, Chronos (ansari2024chronos) discretizes series via bucket quantization, learning more flexible categorical distributions by cross-entropy. Still, discrete tokenization is applied at each time point, which can lead to long contexts. Also, the final performance can be sensitive to quantization techniques. Unlike before, we tame Transformers as native time series foundation models, learning flexible distributions without discrete tokenization.
@@ 2.3 Generative Modeling for Time Series
By addressing complicated distributions during pre-training, generative modeling has become a focal point in the development of various foundation models (zhao2023survey; liu2024sora). While this direction for time series mostly focused on time series generation (tashiro2021csdi) and task-specific forecasters (rasul2021autoregressive; shen2023non; kollovieh2024flow), generative modeling for time series foundation models is hardly explored. With the comparable flexibility in distribution learning as language modeling, diffusion denoising (sohl2015deep) and flow-matching (lipman2022flow) have gained increasing prevalence in continuous-valued modalities (lipman2024flow). Compared with diffusion denoising models, flow-matching provides a simple yet efficient framework. With fewer steps involved in the forward and reverse processes, large models based on flow-matching have shown superior performance in image generation (esser2024scaling).
Despite the connection in value continuity, generating images and future time series are fundamentally different tasks due to the autoregressive property of forecasting. Our proposed TimeFlow Loss is designed for autoregressive models to conduct conditional generation, which is a parameterized loss function (zhang2018unreasonable) for arbitrary distributions and enhances representation learning of foundation models.
@@ 3.1 Flow-Matching
Furthermore, lipman2022flow proved the equivalence of optimizing the Conditional Flow-Matching objective:
@@ 3.2 Generative Models for Probabilistic Forecasting
Given a historical observation $x_{1:t}=\{x_{1},...,x_{t}\}$, the target of time series forecasting is to predict future time series $x_{t+1:t+f}=\{x_{t+1},\dots,x_{t+f}\}$. The task can be generally formulated as $p(x_{t+1:t+f}|\mathbf{h}_{t})$, where $\mathbf{h}_{t}=f_{\phi}(x_{1:t})$ is the learned representation from a deep model $f_{\phi}$. In probabilistic forecasting, explicit optimization objectives are utilized to predict the statistics of future series, e.g., MSE or quantile loss, which have specified $p$ as a prior distribution. While using one parametric density generally fits well on a small amount of data, it can be the major bottleneck for scaling time series foundation models. Inspired by the success of large generative models (rombach2022high; openai2023gpt; esser2024scaling), we introduce generative modeling to realize probabilistic forecasting:
$p_{\theta}(x_{t+1:t+f}|\mathbf{h}_{t})=g_{\theta}\big(f_{\phi}({x_{1:t}})\big).$ (3)
$g_{\theta}$ is a small trainable generative network conditioned on the learned representations of $f_{\phi}$, which is jointly optimized with $f_{\phi}$. While the generative model automatically fits the target distribution, it can sample raw predictions and calculating their statistics for probabilistic forecasting. The aim is conceptually related to conformal prediction (vovk2005algorithmic) but models uncertainty beyond prediction intervals.
@@ 4 Approach
In this work, we conduct a univariate pre-training paradigm, which adopts the S3 format proposed by liutimer to address multivariate data. To mitigate value range discrepancy, we conduct normalization on time series individually per variable. Afterwards, we sample varying-length training samples with the maximum context length of $2880$. As a foundation model, Sundial is required to predict on out-of-distribution series with varied lengths during inference.
@@ 4.1.1 Time Series Tokenization
We adopt stationarization (liu2022non), a non-parametric two-stage instance normalization conducted within each sample, which is initially proposed to mitigate non-stationarity of time series. Here, it helps to address temporal distribution shift and outlier ranges in input series, improving generalizability for zero-shot forecasting.
@@
Given a univariate time series $\mathbf{X}=\{x_{1},\dots,x_{T}\}$, it is divided into patches $\mathbf{x}_{i}=x_{1+(i-1)P:iP}$ with the length of $P$. To address non-divisible length, we pad the input at the beginning and use a binary mask $\mathbf{m}_{i}\in\mathbb{R}^{P}$ for each patch to indicate the padded position. It will lead to $N=\lceil T/P\rceil$ such input tokens. Subsequently, we use a shared MLP $\mathbb{R}^{2P}\mapsto\mathbb{R}^{D}$ to embed all patch tokens:
@@
where $\mathbf{h}_{i}\in\mathbb{R}^{D}$ and $D$ is the dimension of token embedding. Unlike point-level quantization (ansari2024chronos), we reserve original values without discrete quantization. It also reduces the context length (in token) of the Transformer.
@@ 4.1.2 Transformer Backbone
Given $N$ token embeddings $\{\mathbf{h}_{i}\}$, we adopt several crucial adaptations on a decoder-only Transformer to obtain per-token representations aggregated from all previous tokens. First, we adapt Pre-LN (xiong2020layer) to improve pre-training stability. Second, we leverage a causal self-attention mechanism with RoPE (su2024roformer) that introduces the position information of patch tokens. It can be formulated as follows (the layer index is omitted for simplicity):
@@
where $\mathbf{W}_{\mathbf{q}},\mathbf{W}_{\mathbf{k}},\mathbf{W}_{\mathbf{v}}\in\mathbb{R}^{D\times d}$ project token embeddings $\mathbf{H}=\{\mathbf{h}_{i}\}$ into $d$-dimensional queries, keys, and values. $\mathbf{R}_{\Theta,t}\in\mathbb{R}^{d\times d}$ is the rotary matrix with rotation degree $(t\cdot\Theta)$. Lastly, we implement FlashAttention (dao2022flashattention) and KV Cache (pope2023efficiently), since these enhancements for deployment are increasingly emphasized in large foundation models (shoeybi2019megatron; rasley2020deepspeed).
@@ 4.1.3 TimeFlow Loss
Given representations $\{\mathbf{h}_{i}\}$ extracted by the last layer of the Transformer, we aim to generate length-$F$ predictions $\widehat{\mathbf{y}}_{i}=\widehat{x}_{1+iP,F+iP}$ at each position $i$ via our autoregressive model. Motivated by the empirical observation that a larger patch size improves the performance in decoder-only Transformers (das2023decoder) while a small patch size can be more flexible to accommodate data of different frequencies, we adopt multi-patch predictions ($F>P$) for pre-training, which also reduces the steps of autoregressive inference.
@@
The training process involves sampling the noised $\mathbf{y}_{i}^{(t)}$, and jointly input it with $t$. The condition $\mathbf{h}_{i}$ is integrated into the flow-matching network via AdaLN (peebles2023scalable). TimeFlow Loss for autoregressive models is formulated as: